OpenStudy (anonymous):

find the volume of the solid generated by revolving the region bounded by y=x^2 and y=sqrt(x) around the x-axis

OpenStudy (amistre64):

find the volume created when you spin sqrt(x) the subtract the volume you get when you spin x^2 the limits of integration being where the 2 curves meet

OpenStudy (anonymous):

so far I have \[x^2=\sqrt{x}\] \[x^2-\sqrt{x}=0\] \[x^2-x^{1/2}=0\] \[2x-1/2x=0\] \[x(2-1/2)=0\] \[x=0, x=3/2\] is this correct so far?

OpenStudy (amistre64):

not quite; if you can recall the shape of the graphs; these meet at only 2 points; 0^2 = sqrt(0) 1^2 = sqrt(1) are the only options available; so your math is errored someplace

OpenStudy (anonymous):

is I wasnt supposed to integrate when I did

OpenStudy (anonymous):

so x=0 and x=1 are the limits

OpenStudy (amistre64):

yes, 0 to 1 recall the integrating a volume of rotation is adding up areas of circles: pi f(x)^2 \[\int_{0}^{1}pi [x^{1/2}]^2dx - \int_{0}^{1}pi[x^2]^2dx\]

OpenStudy (amistre64):

this cam be simplified; but its the same results create a solid shape and dig out the part thats extra

OpenStudy (anonymous):

The part I'm having trouble with is finding the limits

OpenStudy (amistre64):

x^2 = sqrt(x) ; square the sides x^4 = x x^4 - x = 0 x(x^3 - 1) = 0 x = 0; x = 1

OpenStudy (anonymous):

Ok now I think I understand

OpenStudy (anonymous):

Thanks I am going to try and finish the problem myself.

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thanks